Consider the replacement of the 400 MW San Juan Thermoelectric power plant in Puerto Rico,
which uses heavy fuel oil to generate electricity, with a solar photovoltaic farm. The oil plant has a
heat rate of 10,811 Btu/kWh and a capacity factor of 53%.

a. For installation of south-facing fixed tilt PV arrays, with tilt angle equal to the local latitude in
San Juan, how many acres of collector surface area would be required to replace the generation
from the fuel oil plant? Assume that the photovoltaic cells have an efficiency of 17%.

b. By how many acres can the surface area in part (a) be reduced if single-axis tracking is added
(with the collector tilt angle still equal to the local latitude)?

c. What will be the annual fuel cost savings ($) if the oil power plant is retired, if heavy fuel oil
costs $290 per metric ton and has a heating value of 39,500 Btu/kg?
Answer #1

A.San Juan(puerto rico) latitude and longitude = 18° 27' 58.7988'' N and 66° 6' 20.5956'' W Tilt angle = local latitude in San Juan = 18.46 Data for solar radiation for flat plate collector facing south at a fixed tilt can be taken from WBAN No. 11641 = for year = 5.5 KWh/m2/day = 5.5 KWh/m2/day (1 day =3600h)

Formula used-

Solar radiation at that place* Collector Area*Collector efficiency = Heat rate * Output power*Capacity factor

1 KW= 3,412.14 BTU/h

Heat rate* output power* capacity factor = 10,811*400*10^3*0.53 BTU/h = 671699.2855 KW

Collector Area = 671699.2855*3600/5.5*0.17

                           = 25862218.48 m^2= 6390 acres

C.

Cost of oil = 290$ /10^3 kg

Energy produced= 400*10^3 KW = 1364856 *10^3 BTU/h

Heating value = 39,500 Btu/kg

Annual saving= Energy produced*cost of oil/heating value

                         = $ 10020.46177